The nominal yield on 6-month T-bills is 7%, while default-free Japanese bonds that mature in 6 months have a nominal rate of 4.5%. In the spot exchange market, 1 yen equals $0.01. If interest rate parity holds, what is the 6-month forward exchange rate? Round the answer to five decimal places. Do not round intermediate calculations.
Answer #1

6 month forward Rate =Spot Rate*(1+6 month T-bill rate)/(1+6 month Japanese Bond Rate) =0.01*(1+7%)/(1+4.5%) =0.01024 for 1 Yen